### v4

# Efficient Identification of Critical Transitions via Flow Matching: A Scalable Generative Approach for Many-Body Systems

Abstract
We propose a machine learning framework based on Flow Matching (FM) to identify critical properties in many-body systems efficiently. Using the 2D XY model as a benchmark, we demonstrate that a single network, trained only on configurations from a small ($32\times 32$) lattice at sparse temperature points, effectively generalizes across both temperature and system size. This dual generalization enables two primary applications for large-scale computational physics: (i) a rapid ”train-small, predict-large” strategy to locate phase transition points for significantly larger systems ($128\times 128$) without retraining, facilitating efficient finite-size scaling analysis; and (ii) the fast generation of high-fidelity, decorrelated initial spin configurations for large-scale Monte Carlo simulations, providing a robust starting point that bypasses the long thermalization times of traditional samplers. These capabilities arise from the combination of the Flow Matching framework, which learns stable probability-flow vector fields, and the inductive biases of the U-Net architecture that capture scale-invariant local correlations. Our approach offers a scalable and efficient tool for exploring the thermodynamic limit, serving as both a rapid explorer for phase boundaries and a high-performance initializer for high-precision studies.
preprint: APS/123-QED
I Introduction
Monte Carlo—especially Markov chain Monte Carlo (MCMC)—is one of the most powerful computational tools across three fronts: in condensed matter, it is the workhorse for probing phase transitions and complex systems [Landau_Binder_2005]; in quantum many-body physics, quantum Monte Carlo is among the few reliable non-perturbative methods [Sandvik_2010, RevModPhys.73.33]; and in lattice QCD, Euclidean-lattice Monte Carlo provides the only systematically improvable first-principles non-perturbative approach at low energies [RevModPhys.82.1349, PhysRevLett.49.613]. Algorithmically, MCMC targets the Boltzmann measure via detailed balance and ergodicity (e.g. Metropolis, HMC [HMC], cluster [PhysRevLett_Wolff]), with uncertainties governed by the integrated autocorrelation time [Landau_Binder_2005].
However, MCMC encounters critical slowing down as correlation lengths grow near continuous phase transitions and toward the continuum limit; in lattice QCD, topological modes can freeze, inflating autocorrelations and complicating error estimates [Schaefer_2011]. Fermionic models further face the sign problem, whose generic resolution is NP-hard [PhysRevLett.94.170201]. Progress relies on advanced updates [PhysRevLett_Wolff], tempering/multilevel schemes [Hukushima_1996], and ML-assisted samplers; notably, flow-based generative sampling have shown decorrelated proposals and efficiency gains in lattice field theories [PhysRevD_Flow_based_MCMC, kanwar2024flow_based_sampling_lattice_field, abbott2024practical_app_flows, PhysRevD_NF_MC_sampling, PhysRevD.106.014514].
Figure 1: Schematic diagram of our work. Using Flow Matching with U-Net [ronneberger2015unet] inductive biases, we learn a deterministic probability-flow sampler that—trained only on $L_{0}\times L_{0}$ lattices—simultaneously interpolates across temperatures and extrapolates (dashed lines) to physically consistent configurations of $L_{1}\times L_{1}$ lattices with $L_{1}\gg L_{0}$, revealing scale-invariant rules and enabling a fast and stable exploration of critical phenomena near the critical temperature, $T_{c}$. Here $\chi$ denotes the susceptibility, and the green shadow represents the critical regime.
Early successes with models like Normalizing Flows (NFs) demonstrated the potential to eliminate critical slowing down for a single set of system parameters [PhysRevD_Flow_based_MCMC, hackett2025_flow_based_sampling_multimodal_extendedmode]. However, the inefficiency of these initial approaches in exploring parameter space is a significant limitation. Studying critical phenomena requires simulations across various parameters (e.g., temperatures), but training a separate, computationally expensive model for each point is impractical [PhysRevD_NF_MC_sampling]. A key advancement to address this has been the development of conditional models, which learn the system’s behavior across a continuous parameter range from sparse training data, enabling efficient interpolation and extrapolation [PhysRevD_NF_MC_sampling, Gerdes_2023LQF_equi_flow]. These models are often implemented with either architecturally rigid NFs, which can suffer from training instabilities like mode collapse [hackett2025_flow_based_sampling_multimodal_extendedmode, PhysRevDmode_collapse_for_flow_based], or with Diffusion Models (DM) [zhang2025efficient_unbiased_sampling_boltzmann, Wang2024_Diffusion_models_as_stochastic_quantization, Fukushima_2025_SQ_DM, zhu2024diffusionmodelslatticegauge, zhu2025physics_conditioned_DM], which, while robust, typically rely on a computationally intensive iterative sampling process.
Even as conditional models begin to address the challenge of exploring parameter space, a more fundamental and computationally demanding limitation persists in both traditional MCMC and these new generative approaches: the lack of generalizability across system sizes. Systematic studies of critical phenomena and the thermodynamic limit hinge on a sequence of simulations on progressively larger lattices. Under the current paradigm, each lattice size requires an entirely new, independent simulation or model training, a process whose computational cost scales prohibitively. Therefore, a truly efficient framework would not only interpolate across parameters like temperature but also capture the local, scale-independent correlation patterns, enabling extrapolation from a single, small-system training process to much larger lattices. Our work is designed to address this critical gap directly.
In this work, we employ Flow Matching (FM)  [lipman2023flowmatchinggenerativemodeling], a next-generation framework that combines the advantages of these preceding approaches while mitigating their drawbacks. FM directly learns the deterministic vector field of a continuous transformation, known as the probability flow ODE [song2021score_matching]. This provides a more stable regression-based training objective than traditional NFs and enables significantly faster sampling than the stochastic processes of DMs [ddpm, song2021score_matching]. This training stability, architectural flexibility, and sampling efficiency combination makes FM an ideal candidate for building a robust and generalizable conditional sampler. When integrated with the strong inductive biases of a U-Net architecture [ronneberger2015unet], we demonstrate that this framework offers an effective solution to a crucial challenge: achieving a robust dual generalization. We show that a single model, trained only on data from a small lattice ($32\times 32$), can simultaneously interpolate across a continuous temperature range and extrapolate to predict configurations that reflect the essential critical behavior of substantially larger systems (up to $128\times 128$). This capability to generalize local interactions from limited data establishes our approach as a computationally efficient strategy for locating phase transitions in the thermodynamic limit.
The paper is organized as follows. We begin by introducing the theoretical framework of Flow Matching in Sec. II and our simulation setup in Sec. III. Our main numerical results and analyses, covering the model’s dual generalization, performance, and efficiency, are presented in Secs. IV to VII. We then provide a physical interpretation for these capabilities in Sec. VIII. The paper concludes with a discussion of our findings in Sec. IX and an outlook for future work in Sec. X. Further implementation details, more comprehensive results, and the source code for this study are included in the Appendices.
II Theoretical Structure of Flow Matching
Flow Matching bridges continuous NFs and diffusion models by directly learning vector fields (flows) along predefined conditional probability paths between a simple base distribution (e.g., Gaussian noise) and the target data distribution. It trains models by regressing on these conditional velocities rather than relying on stochastic simulations, resulting in stable, scalable training and efficient sample generation via off-the-shelf ODE solvers.
Figure 2: A naive Flow Matching example. When Flow Matching is trained as a deterministic ODE, trajectories from $\mathbf{v_{\theta}}$ deviate from the straight interpolation in Eq. (1). This is because $\mathbf{v_{\theta}}$ learns the expectation of the conditional velocity, $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$, induced by the random pairings of $(\mathbf{x}_{0},\mathbf{x}_{1})$ during training (left panel). Consequently, the first ODE iteration step (middle panel) points toward the target mean; with a finite number of iterations, inference preferentially produces samples near this mean, thereby underrepresenting the target variance and fluctuations.
II.1 Probability Flow ODE
Flow Matching admits arbitrary paths, granting flexibility to optimize sampling speed or modeling capacity [lipman2023flowmatchinggenerativemodeling]. In this work, we adopt a simple linear interpolation between base and data for simplicity. For a target data sample $\mathbf{x}_{1}\sim\mathcal{P}_{data}$ and a standard Gaussian noise sample $\mathbf{x}_{0}\sim\mathcal{N}(0,\mathbb{I})$ of the same dimensionality, we define the path $\{\mathbf{x}_{t}\}_{t=0}^{1}$ , where $t$ is a virtual time parameter indexing the continuous evolution of the distribution. The path at each subscripted time $t$ is given by
$\mathbf{x}_{t}=t\;\mathbf{x}_{1}+(1-t)\;\mathbf{x}_{0}$ (1)
For a fixed coupling $(\mathbf{x}_{0},\mathbf{x}_{1})$, the conditional velocity field is
$\mathbf{v_{t}}(\mathbf{x}_{t}|\mathbf{x}_{1})=\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{1}-\mathbf{x}_{0}$ (2)
But since $(\mathbf{x}_{0},\mathbf{x}_{1})$ are random, $\mathbf{v_{t}}(\mathbf{x}_{t}|\mathbf{x}_{1})$ is also random. The true velocity field $\mathbf{u_{t}}(x)$ is an average of multiple conditional velocity fields, To know this average, consider a distribution $p_{t}(x)$ which satisfies the following boundary conditions.
$p_{t=1}(\mathbf{x}_{t=1})=\mathcal{P}_{data}\quad and\quad p_{t=0}(\mathbf{x}_{t=0})=\mathcal{N}(0,\mathbb{I}).$ (3)
For simplicity, we can denote the configuration of all conditions as the variable $\phi$, which satisfies the distribution $q(\phi)$. The distribution of the probability path $p_{t}(\mathbf{x}_{t})$ can be written as
$p_{t}(\mathbf{x}_{t})=\int\left[\mathcal{D}\phi\right]\;p_{t}(\mathbf{x}_{t}|\phi)\;q(\phi)$ (4)
Here, $p_{t}(\mathbf{x}_{t}|\phi)$ is the conditional probability path. For example, suppose that $\mathcal{P}_{data}$ is the Boltzmann distribution $exp(-\beta H)/Z$. The configuration $\phi$ of conditions is only the spin configurations $\mathbf{x}_{1}$ (generally, the condition can include temperatures, coupling constants, and other information that may affect the target distribution). Consider the path $\mathbf{x}_{t}$ in Eq. (1), because we choose $\mathbf{x}_{0}$ that satisfies the standard Gaussian distribution, therefore, conditional probability $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$ is the Gaussian distribution with mean $t\mathbf{x}_{1}$ and variance $(1-t)^{2}$ [lipman2024flow_matching_guide_code]. The $p_{t}(\mathbf{x}_{t})$ hence reads as
$p_{t}(\mathbf{x}_{t})=\int\left[\mathcal{D}\mathbf{x}_{1}\right]\;\mathcal{N}\left(\mathbf{x}_{t};t\mathbf{x}_{1},(1-t)^{2}\mathbb{I}\right)\;\frac{e^{-\beta H(\mathbf{x}_{1})}}{Z}$ (5)
We can easily check that Eq. (5) satisfies the requirements in Eq. (3).
With conditional velocity field $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$ and probability density $p_{t}(\mathbf{x}_{t})$ in Eq. (4), the continuity equation gives
$\frac{\partial\;p_{t}(\mathbf{x}_{t})}{\partial t}=-\nabla\cdot\int\left[\mathcal{D}\phi\right]\;\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)p_{t}(\mathbf{x}_{t}|\phi)\;q(\phi)$ (6)
Thus, with Bayes’ theorem, the velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$ is
$\mathbf{u_{t}}(\mathbf{x}_{t})$ $=\int\left[\mathcal{D}\phi\right]\;\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)\frac{p_{t}(\mathbf{x}_{t}|\phi)\;q(\phi)}{p_{t}(\mathbf{x}_{t})}$
$=\int\left[\mathcal{D}\phi\right]\;\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)\;p_{t}(\phi|\mathbf{x}_{t})$ (7)
The corresponding ODE to $\mathbf{u_{t}}(\mathbf{x}_{t})$ is also called the probability flow ODE [song2021score_matching]:
$\frac{d\mathbf{x}_{t}}{dt}=\mathbf{u_{t}}(\mathbf{x}_{t})$ (8)
And we want the neural network $\mathbf{v_{\theta}}$ with learnable parameters $\theta$ to learn this velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$. There are several methods for using neural networks to predict velocity fields. One reliable method is to let the neural network $\mathbf{net_{\theta}}$ predict the target sample and calculate the corresponding velocity field at different time steps [JIT]:
$\mathbf{v_{\theta}}(\mathbf{x}_{t},t)\equiv\frac{\mathbf{net_{\theta}}(\mathbf{x}_{t},t)-\mathbf{x}_{t}}{1-t}$ (9)
Which can be derived from Eq. (1) and Eq. (2).
II.2 Loss Function
With the probability flow ODE in Eq (8), the Flow Matching loss $\mathcal{L}_{FM}(\theta)$ can be defined as
$\mathcal{L}_{FM}(\theta)\equiv\int_{0}^{1}dt\int\left[\mathcal{D}\mathbf{x}_{t}\right]p_{t}(\mathbf{x}_{t})\left[\mathbf{v_{\theta}}(\mathbf{x}_{t},t)-\mathbf{u_{t}}(\mathbf{x}_{t})\right]^{2}$ (10)
However, we do not know the exact velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$ from the training data. Thus, use the trick that the gradient of the loss function $\mathcal{L}(\theta)$ to update the neural network $\mathbf{v_{\theta}}$ is only depend on learnable parameters $\theta$, it is equivalent [lipman2023flowmatchinggenerativemodeling] to define conditional Flow Matching loss $\mathcal{L}_{CFM}(\theta)$ that we want to minimize as
$\mathcal{L}_{CFM}(\theta)\equiv$
$\int_{0}^{1}dt\int\left[\mathcal{D}\mathbf{x}_{0}\mathcal{D}\phi\right]p_{t}(\mathbf{x}_{t}|\phi)q(\phi)\left[\mathbf{v_{\theta}}(\mathbf{x}_{t},t|\phi)-\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)\right]^{2}$ (11)
The loss function is similar to the action of the noninteracting fluid system with integration in the entire condition configuration $\phi$ and the initial Gaussian noise. For the path in Eq. (1), conditional velocity field $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$ here is just $(\mathbf{x}_{1}-\mathbf{x}_{0})$ and see the Appendix A for more details.
II.3 Flow of Random Sampling
Although the path in Eq. (1) is straight, but the training paths are created by randomly pairing data $\mathbf{x}_{1}$ and noise $\mathbf{x}_{0}$ with different conditions. Therefore, different conditional trajectories passing through the same $x_{t}$ at time $t$ generally have different directions, and the velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$ is not identical to the constant vector $(\mathbf{x}_{1}-\mathbf{x}_{0})$ (see Fig. 2). What the neural network $\mathbf{v_{\theta}}(\mathbf{x}_{t},t|\phi)$ learned is an expectation over conditional velocities and there is inherent uncertainty about the direction. One way to understand this is to view the training path as a quantum process, where a sample from the initial distribution $\mathbf{x}_{0}$ has the opportunity to reach any target distribution $\mathbf{x}_{1}$. Flow matching allows the neural network to learn the generators corresponding to the path that contributes most, the classical path. Due to this reason, instead of collapsing sampling to a single jump from noise to data, we must integrate the learned velocity field over a few time steps (however, research on the one-step sampling is rapidly expanding [geng2025meanflowsonestepgenerative, frans2025stepdiffusionshortcutmodels]). Concretely, we divide the interval $t\in[0,1]$ into $N$ steps $\{t_{k}\}^{N}_{k=0}$ and numerically solve the problem (see Fig. 2).
$\frac{d\mathbf{x}_{t}}{dt}=\mathbf{v_{\theta}}(\mathbf{x}_{t},t|\phi)$ (12)
using a method such as explicit Euler or a higher‐order Runge–Kutta. Then, we can transport the samples from the distribution $p_{0}$ to $p_{1}$. For more training and sampling details, see the Appendix B.
III Generative Framework and Observable Estimation
III.1 Parameter-Wide Sampling via a Single Network
One of the main advantages of flow-based sampling is that it can generate target samples in parallel and independently, thus avoiding the autocorrelation problem of traditional MCMC sampling. This advantage can be easily maximized in the Flow Matching training framework. To construct the conditional velocity field in Eq. (2), only two types of information are required as conditions ($\phi$): one is the samples ($\mathbf{x}_{1}$) that obey the target distribution, and the other are the parameters (e.g., temperature in this work) corresponding to the samples. The Flow Matching framework can construct probability flows in completely parallel across the above two data dimensions. For example, the target samples $\mathbf{x}_{1}$ in this work have two indices: temperature and ensemble. Assuming that we have data corresponding to $N_{T}$ temperatures within a range, and each temperature has $N_{cf}$ spin configurations, the target sample $\mathbf{x}_{1}$, as a data array, has the shape ($N_{T}$, $N_{cf}$, $L$, $L$), where $L$ is the length and width of the lattice. Accordingly, the trained network $\mathbf{v_{\theta}}(\mathbf{x}_{t},t,T)$ can generate, in parallel, arbitrarily large ensembles of independent $L\times L$ samples across the training temperature range.
III.2 2D XY Model and Physical Observables
To benchmark our framework, we use the standard two-dimensional (2D) XY model, defined on a square lattice of size $N=L\times L$. The system is described by the Hamiltonian:
$H=-\sum_{\langle i,j\rangle}\cos(\Theta_{i}-\Theta_{j})$ (13)
where $\Theta_{i}\in(-\pi,\pi]$ is the angle of the spin at site $i$, and the sum runs over all nearest-neighbor pairs.
The canonical average of any thermodynamic observable $O$ is calculated over an ensemble of configurations:
$\langle O\rangle=\frac{1}{N_{\text{cf}}}\sum_{k=1}^{N_{\text{cf}}}O_{k}$ (14)
where $O_{k}$ is the value of the observable for the $k$-th configuration and $N_{\text{cf}}$ is the total number of configurations in the ensemble. All intensive quantities are averaged over the $N$ lattice sites and the ensemble, as defined in Eq. (14). Throughout this work, the temperature TT is expressed in units where the Boltzmann constant $k_{\mathrm{B}}=1$.
The energy per site is derived directly from the Hamiltonian:
$E\equiv\frac{1}{N}\langle H\rangle$ (15)
The Cartesian components of the total magnetization are:
$M_{x}\equiv\sum_{i=1}^{N}\cos(\Theta_{i}),\quad M_{y}\equiv\sum_{i=1}^{N}\sin(\Theta_{i})$ (16)
and the root-mean-square (RMS) magnetization per site is:
$m\equiv\frac{1}{N}\left\langle\sqrt{M_{x}^{2}+M_{y}^{2}}\right\rangle$ (17)
The zero-field susceptibility is given by the fluctuations of the magnetization:
$\chi\equiv\frac{1}{NT}\left(\left\langle M_{x}^{2}+M_{y}^{2}\right\rangle-\left\langle\sqrt{M_{x}^{2}+M_{y}^{2}}\right\rangle^{2}\right)$ (18)
To define the spin stiffness, we consider the negative Hamiltonian and the spin current restricted to bonds crossing the lattice in the $y$-direction:
$h_{y}\equiv\sum_{\langle i,j\rangle_{y}}\cos(\Theta_{i}-\Theta_{j}),\quad I_{y}\equiv\sum_{\langle i,j\rangle_{y}}\sin(\Theta_{i}-\Theta_{j})$ (19)
The spin stiffness is then defined as [Sandvik_2010]:
$\rho_{s}\equiv\frac{1}{N}\left(\langle h_{y}\rangle-\frac{1}{T}\langle I_{y}^{2}\rangle\right)$ (20)
Finally, the vortex density is obtained by summing the absolute winding number over each elementary $1\times 1$ plaquette ($\square$):
$\rho_{v}\equiv\frac{1}{N}\left\langle\sum_{\square}\left|\frac{1}{2\pi}\sum_{\langle i,j\rangle\in\partial\square}[\Theta_{j}-\Theta_{i}]_{2\pi}\right|\right\rangle$ (21)
where the inner sum is over the four bonds forming the plaquette’s boundary $\partial\square$, and $\cdots]_{2\pi}$ wraps the angle difference into the interval $(-\pi,\pi]$.
The training data, consisting of spin configurations $\mathbf{x}_{1}$, uses angles mapped to the range $0,2\pi)$. The neural network is trained to directly output continuous angular values without an explicit modular arithmetic operation (i.e., no wrapping into $(-\pi,\pi]$ or $0,2\pi)$). This is a valid approach because all relevant physical observables, such as the energy and magnetization components, are calculated using trigonometric functions (e.g., $\cos(\Theta_{i}-\Theta_{j})$ and $\sin(\Theta_{i})$), which are inherently periodic with a period of $2\pi$. The full implementation of our model and the source code for all simulations are made publicly available to ensure reproducibility (see Sec. XI).
Figure 3: Thermodynamic observables for the $32\times 32$ XY model, demonstrating the Flow Matching (FM) model’s ability to increase sampling density via interpolation. The panels show (a) energy per site $E(T)$ and magnetization $m(T)$, and (b) spin stiffness $\rho_{s}(T)$ and vortex density $\rho_{v}(T)$. The sparse points represent MCMC (MC) data, which serves as the training set, sampled at a coarse temperature interval of $\Delta T=0.10$. Each of these points was generated from an ensemble of $5000$ configurations. The denser set of points shows the results from FM, evaluated on a tenfold denser grid ($\Delta T=0.01$), confirming agreement with and smooth interpolation between the training points.
Figure 4: Susceptibility $\chi(T)$ for a $32\times 32$ XY lattice. The points show results from MCMC (MC) simulations, which serve as both the ground truth and the training data for our Flow Matching (FM) model. The three panels demonstrate the model’s performance when trained on this MCMC data at varying levels of precision: (a) a dense grid with temperature spacing $\Delta T=0.05$, (b) $\Delta T=0.10$, and (c) a sparse grid with $\Delta T=0.50$. The training data consists of $5000$ configurations per temperature point. In all cases, the FM model is evaluated on a much finer grid ($\Delta T=0.01$), showing it successfully reproduces the BKT-related features even when trained on the sparsest data set (In (c), Monte Carlo training data is available only at the four temperatures outside the phase transition point.). Since our method directly generates spin configurations rather than predicting physical quantities, it is well-suited for generating high-quality initial spin configurations for Monte Carlo simulations that require high accuracy.
IV Result I: Dense Estimation of Observables via Interpolation
Flow Matching can learn the transfer of the data background distribution through limited data, so the most intuitive benefit is that one can train the model through discrete samples in parameter space and then estimate observables continuously. Just like when one pulls the tablecloth, the petals on the tablecloth will move accordingly. When the tablecloth pulling method is learned and fixed, no matter how many petals are sprinkled, there will be the same movement pattern. Here we show that Flow Matching trained with MC data at intervals of 0.1 in the temperature range of $0.1\sim 2.0$ can be easily sampled with ten times the temperature density (intervals of 0.01). Figs. 3 show that the spin configuration sampling results of Flow Matching still retain good physical observable trends after increasing the temperature density.
Flow Matching provides an additional benefit and advantage in parameter interpolation. Certain physical quantities, such as susceptibility, often exhibit complex and abrupt changes near phase transition points as temperature varies. When data points are sparse along the temperature axis, conventional numerical fitting or machine learning methods typically struggle to perform effective interpolation. However, although these physical observables change abruptly and complexly, the underlying probability distributions of the samples may vary simply and uniformly concerning certain parameters, such as temperature. In the example of the two-dimensional XY model, as long as Flow Matching accurately learns the relationship between the probability distribution of samples and temperature, it can effectively reconstruct the trend of observables within temperature ranges where training data are sparse. Fig. 4 shows that Flow Matching, trained on spin configuration samples that are extremely sparse in temperature, can still produce configurations that preserve the essential features of the phase transition.
V Result II: Extrapolation for Finite-Size Scaling Analysis
Figure 5: Size extrapolation results for the Flow Matching (FM) model. The model was trained on data from a $32\times 32$ system ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T) and then used to generate configurations for a $64\times 64$ lattice. The generated observables—(a) magnetization, (b) susceptibility, (c) spin stiffness, and (d) vortex density—are shown to be in good agreement with reference MCMC results for the larger system.
A key strength of our Flow Matching approach is its ability to extrapolate to different system sizes without retraining. This capability is central to our study’s primary goal: to demonstrate that Flow Matching can efficiently capture the critical scaling behavior of a system from limited data, prioritizing the identification of phase boundaries over precise quantitative reproduction. As shown in Fig. 5, a single model trained exclusively on $32\times 32$ configurations successfully generates samples for a $64\times 64$ system. The results for key observables exhibit consistent trends with the reference MC data, confirming that the learned local interaction patterns are generalized effectively.
Figure 6: Size extrapolation of magnetization using a single Flow Matching (FM) model trained only on a $32\times 32$ lattice ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The model generates configurations for various system sizes ($L=16,64,128$), and the resulting magnetization is compared against MCMC (MC) data. The FM results match the reference data for each individual size and correctly exhibit the characteristic finite-size scaling trend of the transition.
Figure 7: Susceptibility $\chi(T)$ for various system sizes ($L=16,64,128$), demonstrating the size-extrapolation capability of a single Flow Matching (FM) model. The model was trained exclusively on $32\times 32$ data ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). Although quantitative deviations from the MCMC (MC) reference increase as the system size diverges from the training size, the FM model qualitatively captures the essential physics. Specifically, it correctly reproduces the characteristic shift and sharpening of the susceptibility peak with system size, a key signature of the BKT transition. This property allows us to train with only small-sized Monte Carlo data and directly generate high-quality, larger-sized initial spin configurations, which is beneficial for subsequent large-sized Monte Carlo simulations and greatly improves simulation efficiency.
This extrapolation is robust across a wide range of system dimensions. As shown in Figs. 6 and 7, the same model trained on $32\times 32$ generates configurations for lattices as small as $16\times 16$ and as large as $128\times 128$. It is crucial to note that the model was never exposed to the global statistics of the larger systems during training. The emergence of size-dependent features, such as the shifting susceptibility peak, arises because the model applies learned local interaction patterns consistently across the larger lattice geometry. Although quantitative deviations increase with the extrapolation range (likely due to the accumulation of integration errors and the lack of exact long-range constraints in the training data), the model successfully reproduces the essential finite-size effects and the characteristic BKT scaling signatures.
Figure 8: Finite-size scaling of the BKT transition. The susceptibility peak temperature $T_{\text{peak}}(L)$ is plotted against $(\ln L)^{-2}$ for various lattice sizes (L=16, 20, 24, 28, 32, 40, 48, 64, 80, 96, 128L=16,\;20,\;24,\;28,\;32,\;40,\;48,\;64,\;80,\;96,\;128). All data points were generated from a single Flow Matching model trained exclusively on $32\times 32$ data ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The linear trend, when extrapolated to the thermodynamic limit ($L\to\infty$), yields $T_{\text{BKT}}\approx 0.907$, which is very close to the established value of $\approx 0.8935$ [Hsieh_2013_Finite_size_scaling]. This result confirms that the model has successfully captured the underlying scale-invariant physics. Note that it is neither possible nor necessary to plot the Monte Carlo simulation results here for comparison, since we are using Monte Carlo data of a single size for training.
A finite-size scaling analysis tests the model’s ability to learn scale-invariant physics. Our primary objective here is to qualitatively demonstrate this emergent capability, rather than to achieve a high-precision estimate of $T_{BKT}$. We acknowledge that spin stiffness is a more rigorous metric for analyzing the BKT transition [Hsieh_2013_Finite_size_scaling]. However, as discussed in our error analysis (Sec. VI), the model’s predictions for fluctuation-related quantities are inherently less precise, a characteristic compounded by our deliberate use of a limited, low-precision training dataset (1500 configurations per temperature from MCMC). Consequently, the statistical noise in the generated stiffness data, particularly near the critical point, was too significant to permit a reliable scaling analysis. We transparently document this behavior alongside other observables in Appendix C.
In contrast, the location of the susceptibility peak, $T_{\text{peak}}(L)$, provided a more robust and clearly identifiable feature for demonstrating the expected scaling behavior. We therefore adopted this metric for our analysis. Using the single model trained on $32\times 32$ data, we located $T_{\text{peak}}(L)$ for various lattice sizes ($L=16$ to $128$). As depicted in Fig. 8, a plot of $T_{\text{peak}}(L)$ versus $(\ln L)^{-2}$ reveals the expected linear relationship [PhysRevB.65.184405]. Extrapolating to the thermodynamic limit ($L\to\infty$) yields an estimated $T_{\text{BKT}}\approx 0.907$, in reasonable agreement with the established value of $\approx 0.8935$ [Hsieh_2013_Finite_size_scaling]. This result, achieved without high-precision training data or retraining, suggests that the model has successfully learned the scale-invariant correlation patterns necessary for locating the critical point of the XY model, highlighting its efficiency for studying critical phenomena.
VI Result III: Deviation Analysis and Limits of Estimation
To quantitatively assess the performance of our Flow Matching model, we compute the mean squared error (MSE) between its generated estimates and the reference MCMC data for key physical observables, as defined in Eq. (22), where TT is temperature and $N_{T}$ is the number of temperature points. Fig. 9 presents this error analysis for a model trained on $32\times 32$ lattices and then used to generate samples for systems of size L=16, 32, 64,L=16,\;32,\;64, and $128$.
$MSE\equiv\frac{1}{N_{T}}\sum_{T}\left[\langle\mathcal{O}_{FM}(T)\rangle-\langle\mathcal{O}_{MC}(T)\rangle\right]^{2}$ (22)
Figure 9: Mean squared error (MSE) as a function of temperature, quantifying the deviation of the size extrapolation method. The MSE is calculated between the Flow Matching and reference MCMC results for several observables: susceptibility $\chi$, spin stiffness $\rho_{s}$, energy $E$, magnetization $m$, and vortex density $\rho_{v}$. A single FM model, trained on a $32\times 32$ lattice ($T\in[0.1,2.0]$, $\Delta T=0.01$), was used to generate data for systems of size L=16, 32, 64,L=16,\;32,\;64, and $128$. The evaluation for L=16, 32L=16,\;32 and $64$ uses the same temperature grid as the training, while the evaluation for the largest system, $L=128$, is performed on a finer grid focused on the critical region ($T\in[0.8,1.05]$, $\Delta T=0.005$).
The analysis reveals two important trends. First, the deviations for first-moment quantities remain minimal, indicating that the generated configurations retain sufficient structural information even when extrapolating to system sizes four times larger than the training data. The errors are small for the $L=16$ system, demonstrating robust interpolation to smaller sizes, and predictably increase with the extrapolation range to $L=64$ and $L=128$. Second, and more revealingly, the model exhibits higher consistency for first-moment quantities like energy ($E$) and magnetization ($M$) compared to second-moment quantities related to fluctuations, such as susceptibility ($\chi$) and spin stiffness ($\rho_{s}$). This discrepancy is most pronounced near the critical temperature (see Fig. 7), where critical fluctuations dominate.
This behavior stems from several factors inherent to the Flow Matching framework. Primarily, as discussed in Sec. II.3, the model learns to approximate the mean of the conditional velocity field, $\mathbf{u_{t}}(\mathbf{x}_{t})=\mathbb{E}[\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)]$. This averaging process inherently smooths the learned flow. And it is a direct consequence of the bias-variance tradeoff introduced by the mean squared loss function used for training. According to the loss function in Eq. (11), the expected error of the learned velocity field $\mathbf{v_{\theta}}$ can be decomposed as:
$\mathbb{E}[(\mathbf{v_{\theta}}-\mathbf{v_{t}})^{2}]=\left(\mathbb{E}[\mathbf{v_{\theta}}-\mathbf{v_{t}}]\right)^{2}+\text{Var}[\mathbf{v_{\theta}}-\mathbf{v_{t}}]$ (23)
To minimize this total error, the model must balance its variance (due to finite training data) against a systematic bias towards the mean of the true conditional velocity field. This optimization naturally produces a learned vector field that prioritizes the mean field trend over the precise variance of the target distribution. Furthermore, this initial inaccuracy in the learned field’s variance is compounded during the inference stage. Sampling requires numerical integration of the probability flow ODE with a finite number of steps. As suggested in Fig. 2, the vector guiding the trajectory in each discrete step of the numerical solver is an approximation. Minor errors in the velocity field, particularly regarding its variance, accumulate during the iterative integration process. This can cause the final generated configurations to underestimate the actual spread of the target distribution, leading to a larger error in the variance of the resulting ensemble. Consequently, while the generated samples accurately reflect first-moment quantities such as energy and magnetization, the fluctuation-related properties of the generated ensemble are less reliable. Despite FM-generated samples showing larger errors in fluctuation-related physical statistics, which grow more significantly with size extrapolation, the model remains robust in identifying the temperature location of characteristic peaks in susceptibility. This makes it a highly efficient tool for locating phase transitions and performing finite-size scaling analysis, even if the absolute magnitude of fluctuations is underestimated.
In the training size ($32\times 32$), we also analyze the distribution error of the energy versus the different temperature. We use Jensen–Shannon divergence (JSD) as a measure of the difference between Flow Matching and MCMC sampling distributions. For two discrete probability distributions $\mathcal{P}$ and $\mathcal{Q}$ in the same domain $\mathcal{X}$ define the midpoint distribution.
$\mathcal{M}\equiv\tfrac{1}{2}\bigl(\mathcal{P}+\mathcal{Q}\bigr).$ (24)
The Jensen–Shannon divergence (JSD) is given by
$\operatorname{JSD}(\mathcal{P}\|\mathcal{Q})\equiv$
$\tfrac{1}{2}\sum_{x\in\mathcal{X}}\mathcal{P}(x)\log\frac{\mathcal{P}(x)}{\mathcal{M}(x)}+\tfrac{1}{2}\sum_{x\in\mathcal{X}}\mathcal{Q}(x)\log\frac{\mathcal{Q}(x)}{\mathcal{M}(x)}.$ (25)
With natural logarithms, $0\leq\operatorname{JSD}(\mathcal{P}\|\mathcal{Q})\leq\ln 2$, the upper bound $\ln 2$ is reached precisely when $\mathcal{P}$ and $\mathcal{Q}$ are completely unrelated. Fig. 10 shows the JSD of energy distribution at different temperatures. Across most of the temperature range, including the phase transition point ($T\approx 1.0$ for system size $32\times 32$), the JSD is less than 0.1, indicating that the generated energy distributions are statistically consistent with the Monte Carlo reference. Only near zero temperature does the JSD exceed 0.1. The histograms embedded in Fig. 10 show that the energy distribution becomes increasingly concentrated as the temperature decreases (the bar widths in all histograms are the same). Because Flow Matching uses only one neural network model that is simultaneously trained and sampled over the entire temperature range, the figure reveals that under the same conditions, the generative model faces challenges in resolving the fine-grained variance of extremely narrow distributions. The bias of distributions near the phase transition point also indicates that the neural network is still some distance from generating distributions that are entirely consistent with physics. However, it is still sufficient to quickly locate the phase transition point and generate a high-quality initial spin configuration for Monte Carlo simulation.
Figure 10: Jensen–Shannon Divergence (JSD) between the energy distributions generated by the Flow Matching (FM) and MCMC methods as a function of temperature. The analysis was performed on a $32\times 32$ lattice, using an FM model trained and evaluated under the specified conditions ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The insets show the corresponding energy histograms at four representative temperatures. All histograms are plotted with identical bin widths to compare the distributions’ spread visually. As can be seen in the figure, JSD has a significant value near zero temperature ($T=0.1$) and the phase transition point ($T\approx 0.89$), which means that the neural network still cannot perform perfectly in the extreme distribution region, but it is sufficient for quickly locating the phase transition point and generating high-quality initial spin configurations for Monte Carlo simulation.
VII Result IV: Efficiency in Multi-Scale Critical Analysis
The primary computational advantage of our Flow Matching approach stems from its ability to generalize across system sizes, allowing for rapid estimation without re-simulation. We compare the total time cost required for a typical finite-size scaling analysis to quantify the practical implications of this methodological approach. For our FM model, this involves a one-time training cost on a small lattice ($32\times 32$), followed by rapid inference on a series of larger lattices. The MCMC method requires running a separate, full simulation for each lattice size. We use our highly optimized, GPU-accelerated MCMC algorithm (see Appendix F) as the baseline for this comparison. All timing benchmarks were performed on a single workstation; detailed hardware and software specifications are provided in Appendix G. The key metric for this evaluation is the Sampling Cost, defined as the average wall-clock time to generate a single independent configuration:
$\text{Sampling Cost}\equiv\frac{t_{\text{sample}}}{N_{\text{sample}}}$ (26)
where $t_{\text{sample}}$ is the total time required to generate $N_{\text{sample}}$ independent samples across all temperatures and ensembles.
For the FM model, the overall cost has two components: the inference cost, given by Eq. (26), and a one-time, upfront training cost. The average time per configuration during the training phase is defined as:
$\text{Training Cost}\equiv\frac{t_{\text{train}}}{N_{\text{train}}}$ (27)
where $t_{\text{train}}$ is the total time to process $N_{\text{train}}$ training configurations. The training cost here is the amortized time over the total training samples.
Figure 11: Computational cost per independent configuration as a function of system size $L$. The figure compares the performance of our GPU-accelerated MCMC method against the Flow Matching (FM) model, which was trained only on the $L=32$ system. The FM cost is presented in two ways: the inference cost alone and the total cost including the amortized one-time training expense, as defined in Eq. (27). The results show a crossover in efficiency, which can be used to locate phase transition points across multiple sizes quickly or to rapidly generate high-quality initial spin configurations for large-scale Monte Carlo simulations. Because the cost of training is a one-time expense. This method reduces the burden of traditional Monte Carlo simulations by leveraging the ability of trained neural networks to be reused or fine-tuned for other tasks.
As shown in Fig. 11, while the initial training constitutes a significant upfront investment, this cost is amortized over all subsequent sample generation for various larger system sizes. The crossover point at $L=128$ demonstrates that our method offers a streamlined workflow for locating critical points in the thermodynamic limit. This highlights that the actual efficiency of our model lies not in single-sample speed, but in its ability to circumvent the need for repeated, costly simulations across different scales.
VIII Physical Interpretation of Size Extrapolation
The model’s remarkable ability to generalize across system sizes stems from a carefully engineered synergy between the learning framework, the model architecture, and the fundamental nature of the physical system. Unlike methods that learn size-specific global statistics, our approach is designed to distill the underlying, scale-invariant local correlation patterns dictated by the Hamiltonian. The following subsections will elaborate on the key components that enable this capability. We will first discuss how the strong inductive biases of Convolutional Neural Networks (CNNs), namely locality and translation equivariance, are naturally aligned with the structure of local physical interactions. We will then explain how the specific U-Net architecture leverages these biases across multiple scales to capture both local details and emergent long-range correlations. Finally, we will synthesize these points to argue that the model learns a representation of the local generative dynamics for locating critical points.
VIII.1 Inductive Bias
Inductive bias refers to implicit assumptions within machine learning algorithms that enable generalization to unseen data. The effectiveness of an algorithm depends significantly on how well its inductive biases align with the characteristics of the given task. Thus, selecting an appropriate model involves matching these biases to the specific nature of the problem. For example, several recent studies similar to this work have pointed out that the equivariance of neural networks (a kind of inductive bias) has significant advantages in dealing with physical systems with symmetries [Schuh:2025fky_Hubbard_Model_NF, Equivariant_flow_matching].
Convolutional Neural Networks (CNNs) inherently contain strong inductive biases that are advantageous for processing natural images whose statistical properties align closely with these biases. Consequently, CNNs efficiently learn visual representations with minimal data. These biases are equally effective in physical systems characterized by local interactions. The core computational mechanism of CNNs is convolution, typically implemented using small convolutional kernels (e.g. $3\times 3$ in this work). This design explicitly encodes the inductive bias of locality, which posits that the most relevant information for interpreting any given part of an image is predominantly contained within its spatial vicinity. Thus, the significance of any pixel largely depends on its neighboring pixels, allowing CNN neurons to focus solely on small, localized receptive fields. This notion parallels the physical treatment of complex many-body systems, where interactions are often defined locally. Physicists commonly represent such systems using local Hamiltonians, and physical fields or spin configurations governed by these Hamiltonians naturally conform to CNN’s locality inductive bias, where the state or behavior at a given location is determined by its neighbors.
A second crucial inductive bias inherent in CNNs is translation equivariance, arising from parameter sharing across convolutional layers. Within each convolutional layer, the same kernel systematically traverses the entire input feature map, ensuring that the features and behaviors learned at any specific location can directly inform the interpretations at all other spatial locations. This bias aligns with the practice in physics of using uniform or similarly structured local Hamiltonians to describe complex many-body systems consistently throughout their spatial domain, rather than employing entirely distinct physical mechanisms or Hamiltonians at different locations.
The powerful inductive biases embedded in CNN architectures, particularly locality and translation equivariance, are pivotal to their efficiency in image processing tasks. They also intrinsically resonate with the fundamental properties and representations of physical many-body systems when dealing with the data at the physical field or spin configuration level.
VIII.2 U-Net Structure
While the inductive biases of locality and translation equivariance make standard CNNs well suited to systems with local interactions, they struggle with long-range correlations near criticality because the receptive field grows only linearly with depth of network. The U-Net architecture [ronneberger2015unet], built by the encoder and decoder structure, mitigates this by efficiently integrating information across multiple spatial scales, especially when paired with information-preserving resizing (e.g., pixel-shuffle–based rearrangements (see Fig. 12[shi2016pixel_shuffle]). Crucially, even when operating on downsampled feature maps in the encoder that summarize large-scale structure, convolution remains a local operator: kernels still act on neighboring sites of the current lattice, thereby learning transformation rules that are consistent across resolutions—i.e., effective local interactions in renormalized coordinates that remain (approximately) scale-invariant. To ensure that the model accurately reflects the underlying physics, we align its architectural choices with the simulation’s boundary conditions. All Monte Carlo simulations in this work employ periodic boundary conditions. Correspondingly, all convolutional layers within the U-Net architecture are implemented with circular padding. This crucial detail enforces the same periodic topology on the feature maps as on the physical lattice, ensuring that the learned local rules are consistent with the system’s global structure.
Figure 12: The pixel shuffle operation [shi2016pixel_shuffle], a lossless resizing technique used in our U-Net architecture. The encoder employs a space-to-depth pixel shuffle (illustrated from left to right) to downsample spatial resolution while increasing channel depth. Conversely, the decoder uses an inverse depth-to-space pixel shuffle (right to left) for upsampling to reconstruct the spatial dimensions.
Mechanistically, the encoder uses an pixel shuffle (space-to-depth) [shi2016pixel_shuffle] to reduce spatial resolution without discarding information, mapping, for example, each $2\times 2$ block into four channels (see Fig. 12). This preserves translation equivariance and expands the effective receptive field exponentially with depth [wang2024theoreticalanalysis_inductive_biases], while convolutions at coarser resolutions still implement local updates on block variables. The decoder applies the inverse pixel shuffle (depth-to-space) to reconstruct high-resolution outputs (see Fig. 12), and skip connections concatenate encoder features with aligned decoder features. This fusion preserves fine-grained locality while enforcing cross-scale consistency, allowing the network to integrate global context without sacrificing the learning of scale-consistent local rules at each resolution. For more architectural and technical details, see the Appendix D.
For many-body physical systems, this design naturally aligns with a renormalization-group perspective: space-to-depth provides a lossless coarse-graining that exposes emergent long-range correlations at the bottleneck, while convolutions at every level learn local interaction laws on the corresponding (reindexed) lattice. Skip connections then reinsert microscopic details, ensuring that global, system-spanning patterns inform but do not overwrite local physics. As a result, U-Net can recognize multi-scale signatures of criticality while maintaining the learning of (approximately) scale-invariant local physical rules.
To empirically validate that the U-Net’s inductive biases are key to its successful size extrapolation, we conducted a comparative study against a Diffusion Transformer (DiT) model [peebles2023DiT], which lacks strong, built-in locality. The U-Net significantly outperformed the DiT, especially in the extrapolation task, despite having substantially fewer parameters. This result supports our claim that architectural alignment with the problem’s local nature is critical. The detailed experimental setup, results, and discussion of this comparison are presented in Appendix E.
VIII.3 Capturing Local Dynamics for Scalable Generation
The MCMC simulation is inherently tied to the global parameters of the system it simulates. It learns nothing beyond the equilibrium properties of the size and temperature of that single system. The process is a powerful tool for the analysis of a fixed system, not for learning generalizable interaction patterns across scales. In contrast, Flow Matching operates on a different principle. It does not sample a stationary distribution but instead learns a deterministic, non-equilibrium transformation. The goal is to construct a continuous path that transforms a simple, unstructured probability distribution (e.g., Gaussian noise) into the complex target data distribution (in our case, the physical spin configurations from MCMC). The characteristics of this generative framework include:
  1. 1.
    Directed Dynamics : The process is governed by an ODE defined by a learned vector field, This vector field acts as a time-dependent force that deterministically ”flows” any initial random state $\mathbf{x}_{0}$ at time $t=0$ to a structured, physical state $\mathbf{x}_{1}$ that is statistically consistent with the physical ensemble. at time $t=1$. Unlike MCMC, this process is explicitly designed to change the probability distribution over time, moving it along a predefined path from noise to data.
  2. 2.
    Approximating Local Dynamics of Transformation : The model is trained to minimize the discrepancy between its vector field and a target vector field derived from the MCMC data. Crucially, the model is not learning the global statistical properties of the small training systems. Instead, it is learning a function that dictates, for any given configuration $\mathbf{x}_{t}$ at any intermediate time $t$, the optimal infinitesimal change required to align the configuration with the target distribution. For the XY model, this involves learning local update tendencies, such as encouraging adjacent spins to align to lower the energy or resolving topological defects like vortices.
  3. 3.
    Distillation of Scale-Invariant Patterns: The Hamiltonian of the XY model is defined by local, nearest-neighbor interactions. The physical principles governing low-energy configurations, the preference for smooth spin variations and the structure of vortices, are therefore local and independent of the overall system size. Flow Matching, by learning the vector field from local configuration data, distills these scale-invariant interaction patterns. Powered by the inductive biases embedded in CNN architectures, the learned vector field at specific site effectively only ”sees” the configuration of its local neighborhood, not the global lattice boundaries.
The origin of system-size generalization lies in the distinction between learning global statistics and approximating local dynamics. Flow Matching, in our framework, captures the local conditional velocity field—effectively, the ”force” required to align spins locally based on their immediate environment. Because these local dynamics are translationally invariant and independent of the total system volume, they can be seamlessly applied to larger lattices.
Crucially, this explains how finite-size effects (FSE) are captured despite the model being size-agnostic. FSEs are not explicitly learned or memorized from the small training lattice. Instead, they emerge during the inference process as a consequence of applying the learned local dynamics on a manifold with larger spatial dimensions and periodic boundary conditions. The global correlation length is allowed to grow naturally until it is bounded by the new system size ($L$), thereby reproducing the correct scaling behavior ($T_{peak}(L)$) without requiring size-specific training data.
Consequently, applying the trained model to a larger lattice does not extrapolate global statistical properties; it simply deploys the same learned local rules over a broader domain. The model generates valid large-scale configurations because the underlying physics it has internalized is the same at all scales. In essence, it learned to approximate the local generative process of the XY model, thereby enabling the construction of physically plausible configurations of arbitrary size. This success underscores a powerful principle in physics-informed machine learning: the most effective models for locating phase transitions are often those whose architectural inductive biases are deliberately aligned with the intrinsic locality and symmetries of the physical system being studied, a principle that has proven effective in a growing body of literature at the intersection of machine learning and physics [Gerdes_2023LQF_equi_flow, Schuh:2025fky_Hubbard_Model_NF, Equivariant_flow_matching, PhysRevLett.125.121601, PhysRevD.103.074504, zhu2025physics_conditioned_DM, Hashimoto_2024].
IX discussion
IX.1 Comparison of Normalized Flow and Flow Matching
Both Normalizing Flows (NFs) and Flow Matching (FM) are powerful generative frameworks designed to learn a transformation between a simple, tractable base distribution and a complex target data distribution. However, they differ fundamentally in their approach, leading to significant implications for model architecture, training stability, and flexibility (see Fig. 13).
Normalizing Flows construct this transformation as a sequence of invertible and differentiable mappings. This framework’s core requirement is that each transformation’s Jacobian determinant must be computationally tractable. This is essential because NFs rely on the change of variables formula to compute the target distribution’s exact likelihood, which is typically used to minimize the Kullback-Leibler (KL) divergence during training. This constraint imposes severe restrictions on the neural network architectures that can be employed. To ensure a tractable Jacobian, NFs often utilize specialized structures like coupling layers [real_nvp, nice], which partition the input dimensions into active and passive sets. Consequently, these models require carefully designed problem-specific masking mechanisms to ensure that all dimensions are updated over a sequence of layers. This architectural rigidity makes the implementation more complex and potentially unstable, and may also break the inherent symmetries of the modeled system [PhysRevLett.125.121601, PhysRevD.103.074504], a significant drawback in physical applications.
Flow Matching, on the other hand, circumvents these limitations by reframing the problem. Instead of learning a series of discrete, invertible maps, FM learns the continuous-time vector field—the generator—of a probability flow that transports the base distribution to the target distribution. The model is trained via a straightforward regression objective: minimizing the mean squared error between the network’s predicted vector field and a target vector field derived from simple interpolations between noise and data. Since this process does not require the computation of Jacobian determinants during training, it is free from the rigid architectural constraints that limit NFs. This allows highly flexible and powerful architectures, such as the U-Net employed in this work, without the need for intricate masking schemes or concerns about invertibility. The resulting training process is typically more stable and scalable, highlighting a key advantage of FM in building robust and generalizable generative models. By focusing on learning the flow generator, FM provides a more direct and versatile framework, which we will discuss further.
Figure 13: A schematic comparison of three sampling methodologies. Traditional Markov Chain Monte Carlo (MCMC) operates by randomly sampling configurations from a system’s phase space according to probability weights, a process that can be computationally intensive. Normalizing Flows (NFs) offer a generative alternative by constructing a sequence of invertible maps, but their reliance on a tractable Jacobian determinant imposes significant constraints on the neural network architecture. In contrast, Flow Matching (FM), the approach used in this work, directly learns the vector field (flow) that generates the transformation between distributions. This allows for greater architectural flexibility and a more stable training objective, as samples are transported deterministically along a predefined path.
IX.2 Generator Learning Framework
Learning the generator in Flow Matching methods, specifically the velocity field along the probability path connecting different distributions, provides numerous advantages. By explicitly modeling this generator through a neural network, the mapping between distributions becomes explicit and mathematically tractable, circumventing the implicit representations typically encountered in conventional neural networks. This explicit representation significantly improves the flexibility and applicability of the neural network model. Furthermore, framing the neural network as the generator of the distribution mapping, it becomes feasible to employ diverse training strategies beyond Flow Matching alone, such as traditional maximum likelihood estimation or reverse KL divergence minimization in training Normalizing Flows (but need to solve the probability flow ODE during the training process), which can also be adapted to suit various problem contexts and specific research objectives. Overall, this approach aligns well with contemporary efforts in machine learning to build more interpretable, flexible, and versatile generative models, enabling clearer insight into the learned representations and improving the model’s adaptability across a broader range of applications.
IX.3 Limitations of Flow Matching
A significant limitation highlighted by our analysis is the model’s tendency to underestimate second-moment observables, which are directly related to physical fluctuations. This is evident in the increased MSE for susceptibility as system size grows (see Fig. 9) and, more strikingly, in the statistical noise observed in the extrapolated spin stiffness, as detailed in Appendix C. This behavior is likely rooted in the nature of the regression-based training objective (Eq. 11), which prioritizes the mean-field trend of the conditional velocity field, often at the expense of capturing the full variance. This characteristic, in fact, limits the utility of the generated ensemble for directly extracting critical exponents that rely on precise fluctuation magnitudes. However, as shown in Sec. V, the model remains highly effective for locating the critical temperature ($T_{c}$). Improving the model’s ability to capture the variance of the target distribution faithfully is therefore a critical challenge for future research. While potential avenues could include exploring alternative loss functions, a more immediate and practical solution is integrating our method into a powerful hybrid workflow. In this paradigm, the configurations generated by the FM model serve as high-quality, decorrelated initial states for a subsequent, brief refinement phase using traditional MCMC. A short MCMC chain can efficiently correct for any residual biases from the generative process, ensuring the final ensemble rigorously converges to the target Boltzmann distribution. Combining initial MCMC seeding, massive-scale FM sampling for interpolation and extrapolation, and a final MCMC refinement preserves the core computational advantage. The Flow Matching stage still acts as a potent accelerator, significantly reducing the computational cost required to obtain a large set of independent samples suitable for precision tasks like accurately determining critical exponents.
Another limitation is that the Flow Matching framework is designed for systems with continuous degrees of freedom, as in the XY model. Its extension to discrete models (e.g., Ising or Potts) is non-trivial, as the continuous probability flow is not directly applicable. There are two main paths to address this challenge. The first involves reformulating the discrete problem into an equivalent continuous one, for instance, using an auxiliary field method like the Hubbard-Stratonovich transformation, making the system tractable for our approach. A second, more direct path is to leverage generative models explicitly designed for discrete spaces. The development of frameworks such as Masked Diffusion [masked_diffusion] and more recent Discrete Flow Matching [gat2024discrete_flow_matching] marks a promising research frontier for applying these powerful sampling techniques to a broader class of physical systems. Indeed, several recent works have already started to explore such applications [tuo2025scalablemultitemperaturefreeenergy].
X Conclusion and Outlook
In this work, we have introduced a generative model based on Flow Matching that offers an efficient framework for estimating critical properties in many-body systems. Our key finding is that a single network, by capturing the scale-invariant local interaction patterns of the Hamiltonian rather than the global statistics of a specific system size, can generate configurations that reflect essential critical correlations across continuous temperatures and, crucially, for significantly larger lattices. This dual-generalization capability fundamentally alters the workflow for studying critical phenomena. It provides an efficient ”train-small, predict-large” strategy that can complement the traditional, brute-force approach of running separate, costly simulations for each system size. Our analysis shows that this approach provides a significant net computational advantage for large-scale studies, making the identification of phase transitions via finite-size scaling analysis more accessible.
The principles underpinning this method’s success—aligning the inductive biases of a CNN architecture with the locality inherent in local interaction models suggest its applicability to a broader class of models. The framework is particularly well-suited for systems described by continuous fields on a lattice, opening promising avenues for its application in lattice field theories, such as scalar $\phi^{4}$ theory. Furthermore, its utility could extend to quantum many-body problems, especially those where quantum Monte Carlo methods map the system onto a classical statistical model with continuous variables. While our analysis highlights the tendency to underestimate second-moment observables, this limitation opens a promising avenue for a powerful hybrid workflow. The core computational advantage is preserved by using the configurations generated by our Flow Matching model as high-quality, decorrelated initial states for a brief MCMC refinement. This synergistic approach, where FM acts as a potent accelerator and MCMC ensures rigorous convergence, paves the way for achieving enhanced efficiency in high-precision tasks, such as accurately determining critical exponents. This establishes our methodology as an efficient sampling tool and a foundational component for next-generation, hybrid simulation strategies in statistical mechanics and beyond.
XI Code Availability
The source code for this work’s models, simulations, and analysis is publicly available. The code can be accessed from the GitHub repository at https://github.com/ToelUl/Flow-to-Field. This repository includes the full implementation of the Flow Matching framework, the U-Net architecture, and the GPU-accelerated MCMC simulations. Further details, including hyperparameter settings, random seeds, and instructions for reproducing the figures, are also provided to ensure full reproducibility.
XII Acknowledgment
We thank Chi-Ting Ho, Hsuan-Yu Wu, Po-Chung Chen, and Lingxiao Wang for valuable discussions and suggestions. This work is supported by the National Center for Theoretical Sciences, the Higher Education Sprout Project funded by the Ministry of Science and Technology, and the Ministry of Education in Taiwan. DWW is supported under the grant NSTC 113-2112-M-007-036.
Appendix A Details of Loss Function
We can show that the gradient of the conditional loss in Eq. (11) is identical to that of the original objective in Eq. (10):
$\nabla_{\theta}\mathcal{L}_{FM}$ $=\int_{0}^{1}dt\int\left[\mathcal{D}\mathbf{x}_{t}\right]p_{t}(\mathbf{x}_{t})\nabla_{\theta}\left[\mathbf{v_{\theta}}^{2}-2\mathbf{v_{\theta}}\cdot\mathbf{u_{t}}\right]$
$=\int_{0}^{1}dt\int\left[\mathcal{D}\mathbf{x}_{0}\mathcal{D}\phi\right]\nabla_{\theta}\left[p_{t}(\mathbf{x}_{t}|\phi)q(\phi)\mathbf{v_{\theta}}^{2}\right$.
+ $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\left.+\;2\mathbf{v_{\theta}}\cdot\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)p_{t}(\mathbf{x}_{t}|\phi)\;q(\phi)\right]$
$=\nabla_{\theta}\mathcal{L}_{CFM}$ (28)
However, it is numerically difficult to calculate the loss function $\mathcal{L}(\theta)$ in Eq. (11) directly. Thus, the strategy is to sample few pairs of $(\mathbf{x}_{0},\mathbf{x}_{1})$ and other conditions randomly in each training step, calculate the corresponding $\mathbf{x}_{t}$, and minimize $\left[\mathbf{v_{\theta}}(\mathbf{x}_{t},t|\phi)-\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)\right]^{2}$ by updating the parameters $\theta$. After repeating the above steps until the neural network converges, we can use learned neural network $\mathbf{v_{\theta}}(\mathbf{x}_{t},t|\phi)$ to approximate velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$. Now the loss function is the expectation value that can be written as
$\mathcal{L}_{CFM}(\theta)=$
$\mathbb{E}_{(\mathbf{x}_{1},\phi)\sim\mathcal{P}_{data},\;\mathbf{x}_{0}\sim\mathcal{N}(0,I),\;t\sim p(t)}\left[\mathbf{v_{\theta}}(\mathbf{x}_{t},t|\phi)-(\mathbf{x}_{1}-\mathbf{x}_{0})\right]^{2}$ (29)
Here, the distribution of time $p(t)$ can be chosen as uniform between 0 and 1. But the better choice of $p(t)$ will be discussed in the next section.
Appendix B Training and Sampling Strategy
According to the path defined in Eq. (1), we can define the effective signal-to-noise ratio (SNR) and the logarithm SNR ($\lambda(t)$) as
$SNR\equiv\frac{t^{2}}{(1-t)^{2}}\;\;,\;\;\lambda(t)\equiv\log SNR=2\mathrm{logit}(t)$ (30)
Experience has revealed that Flow Matching will have a relatively large error when the SNR is 1, so sampling time from a logit normal distribution in Eq. (31) with $\mu=0$ and $\sigma=1$ during the training phase would be a good choice.[esser2024SD3]
$p(t,\mu,\sigma)\equiv\frac{1}{\sigma\sqrt{2\pi}}\frac{1}{t(1-t)}\exp\left(-\frac{\left(\mathrm{logit}(t)-\mu\right)^{2}}{2\sigma^{2}}\right)$ (31)
Since the curvature of the probability flow ODE solution is inconsistent at different times (see Fig. 2). By mapping the time steps, reducing the time-step spacing where the curvature is large and enlarging the time step spacing where the curvature is slight can achieve better sampling quality. Since the time is sampled from the logit normal distribution during the training phase, the log-SNR $\lambda(t)$ in Eq. (30) can be reparameterized with a new mean and standard deviation $\lambda(t,\mu,\sigma)$,
$\lambda(t,\mu,\sigma)\equiv 2\,(\mu+\sigma\,\mathrm{logit}(t))\;\;,\;\;\mathrm{logit}(t)\sim\mathcal{N}(0,1)$ (32)
and the corresponding time-mapping relationship can be solved as
$\lambda(t,\mu,\sigma)=\lambda(t^{\prime})\Rightarrow\,t^{\prime}(t,\mu,\sigma)=\frac{e^{\mu}}{e^{\mu}+(\frac{1}{t}-1)^{\sigma}}$ (33)
The advantage of this time schedule is that one can flexibly adjust the mean and standard deviation to adapt to different sampling situations, and it is an effective means in practice.[esser2024SD3][labs2025_flux1_kontext_flowmatching] In this work, the $(\mu,\sigma)$ are set to $(0,1)$ during training phase and (0.3,1)(-0.3,1) during inference phase.
All models in this work were trained using an identical hyperparameter configuration. We employed the AdamW optimizer [loshchilov2019decoupled_weight_decay] with the 1-cycle learning rate policy [smith2019superconvergence], setting the maximum learning rate to $2\times 10^{-4}$. Training was conducted with a batch size of 32, and we used 4 steps of gradient accumulation, resulting in an effective batch size of 128 for each parameter update.
The training and inference process is detailed in Algorithms 1 and 2. The $\mathbf{x}_{1}$ is the spin configuration (spin angles) generated by the MCMC, $\mathbf{x}_{0}$ is the standard Gaussian noise with the same shape of spin configuration, $c_{\text{data}}$ is the corresponding temperature of the spin configuration and $p_{t}(t,\mu,\sigma)$ is the logit normal distribution in Eq. (31). In the inference stage, since the probability flow ODE needs to be solved, the computational accuracy and time cost must be considered at the same time. On the premise that Flow Matching maintains the efficiency advantage over traditional algorithms, the 3rd-order Heun’s (Heun3) method within 10 steps is reliable. All Flow Matching sampling methods in this work use the 5-step Heun3 method to sample spin configurations.
Input: Prior distribution $\mathcal{N}(0,\mathbb{I})$, time distribution $p(t,\mu,\sigma)$, data distribution $p_{\text{data}}$, conditioning data $c_{\text{data}}$, velocity field model $\mathbf{v}_{\theta}$, number of training steps NN, batch size $B$, learning rate $\eta$
Output: Trained model parameters $\theta$
for $n\leftarrow 1$ to NN do
   Sample $(\mathbf{x}_{1},c)\sim(p_{\text{data}}^{B},c_{\text{data}}^{B})$
   Sample $\mathbf{x}_{0}\sim\mathcal{N}^{B}(0,\mathbb{I})$
   Sample $t\sim p^{B}(t,\mu,\sigma)$
 $\mathbf{x}_{t}\leftarrow t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}$ Conditional flow path
 $\mathbf{v}_{\star}\leftarrow\mathbf{x}_{1}-\mathbf{x}_{0}$ Conditional vector field target
 $\mathcal{L}\leftarrow\frac{1}{B}\sum_{i=1}^{B}\bigl\|\mathbf{v}_{\theta}(\mathbf{x}_{t}^{(i)},t^{(i)},c^{(i)})-\mathbf{v}_{\star}^{(i)}\bigr\|_{2}^{2}$
 $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$ Update model parameters
return $\theta$
Algorithm 1 Flow Matching — Training
Input: Prior distribution $\mathcal{N}(0,\mathbb{I})$, velocity field $\mathbf{v}_{\theta}$, number of steps $K$, conditioning $c_{\text{data}}$, time-mapping function $t^{\prime}(\cdot,\mu,\sigma)$
Output: Sample $\mathbf{x}_{K}\approx p_{\text{data}}$
Sample $\mathbf{x}_{0}\leftarrow\mathcal{N}(0,\mathbb{I})$
Generate uniform time steps $\tau_{k}=k/K$ for $k=0,\dots,K$
Map to non-uniform time steps $t_{k}=t^{\prime}(\tau_{k},\mu,\sigma)$ for $k=0,\dots,K$
for $k\leftarrow 0$ to K1K-1 do
 $h\leftarrow t_{k+1}-t_{k}$ Non-uniform step size
 $\mathbf{k}_{1}\leftarrow\mathbf{v}_{\theta}(\mathbf{x}_{k},t_{k},c_{\text{data}})$
 $\mathbf{k}_{2}\leftarrow\mathbf{v}_{\theta}(\mathbf{x}_{k}+\frac{h}{3}\mathbf{k}_{1},t_{k}+\frac{h}{3},c_{\text{data}})$
 $\mathbf{k}_{3}\leftarrow\mathbf{v}_{\theta}(\mathbf{x}_{k}+\frac{2h}{3}\mathbf{k}_{2},t_{k}+\frac{2h}{3},c_{\text{data}})$
 $\mathbf{x}_{k+1}\leftarrow\mathbf{x}_{k}+\frac{h}{4}(\mathbf{k}_{1}+3\mathbf{k}_{3})$
return $\mathbf{x}_{K}$
Algorithm 2 Flow Matching — Inference with Time-Mapping and 3rd-Order Heun’s Method
Appendix C More Physics Observables in Size Extrapolation
To provide a comprehensive view of the model’s extrapolation performance, this appendix presents a detailed comparison for several key observables across system sizes L=16, 32, 64,L=16,\;32,\;64, and $128$. The model was trained only on $L=32$ data with 1500 configurations per temperature from MCMC. The setup is identical to Sec. V. The results distinguish between the model’s robust performance on first-moment quantities (energy in Fig. 14, vortex density in Fig. 15) and its challenges with the second-moment, fluctuation-related quantity (spin stiffness in Fig. 16).
Figure 14: Size extrapolation of energy per site $E(T)$ for various system sizes ($L=16,64,128$). A single Flow Matching (FM) model, trained only on $32\times 32$ data, generates configurations that show excellent agreement with the reference MCMC (MC) data. This highlights the model’s ability to accurately extrapolate first-moment quantities.
Figure 15: Size extrapolation of vortex density $\rho_{v}(T)$ for various system sizes ($L=16,64,128$). Generated by a single Flow Matching (FM) model trained only on $32\times 32$ data, the results align well with the MCMC (MC) reference. This further validates the model’s robust performance on observables not primarily driven by large-scale fluctuations.
Figure 16: Size extrapolation of spin stiffness $\rho_{s}(T)$ for various system sizes ($L=16,64,128$), generated from a single FM model trained on $32\times 32$ data. The statistical noise visible across all curves, including the MCMC (MC) reference, is partially attributable to the modest ensemble size of 1500 configurations per temperature used for both training and comparison. While the FM model captures the qualitative behavior, a rigorous finite-size scaling analysis for the BKT transition requires locating the intersection of the stiffness curve with the universal line $2T/\pi$, according to the Nelson-Kosterlitz criterion [PhysRevLett.39.1201]. As the figure shows, for extrapolated sizes ($L=64,128$), a noticeable deviation between the FM and MC results appears near the critical transition temperature where this intersection is expected. This discrepancy invalidates a scaling analysis based on the intersection point and justifies our choice to use the more robust susceptibility peak (Sec. V).
Appendix D Neural Network Architecture
Our model employs a U-Net architecture, a design renowned for its efficacy in image-to-image tasks due to its hierarchical structure that captures multi-scale features [ronneberger2015unet]. The network is composed of an encoder path, a bottleneck, and a decoder path, with skip connections linking corresponding levels of the encoder and decoder to preserve high-resolution spatial details. To handle the conditional nature of our generation task, the model integrates time and temperature embeddings, which are processed and fused before modulating the network’s behavior at each block.
D.1 Core Building Blocks
The fundamental component of our U-Net is the ”ResnetBlock” (RB). Each block consists of two main sequential operations: a spatial processing layer and a feedforward neural network. The input to each operation is first normalized using Root Mean Square Layer Normalization, a variant of Layer Normalization known for its computational efficiency and performance [zhang2019RMSNorm].
A key feature of the RB (see Fig. 17) is its adaptive conditioning mechanism, inspired by the Adaptive Layer Normalization (adaLN) used in Diffusion Transformers [peebles2023DiT]. The conditional embedding, derived from time and temperature, is passed through a linear layer to generate shift and scale parameters. These parameters modulate the normalized feature maps within the block, allowing the network’s behavior to be dynamically controlled by the input conditions.
While the U-Net’s convolutional layers excel at learning local physical rules due to their inherent inductive biases of locality and translation equivariance, they are less effective at capturing the long-range correlations that emerge near critical points. To address this limitation, we replace the standard convolutional layers in RB with a specialized multi-head self-attention mechanism [vaswani2023attention_need] in the network’s deeper, lower resolution levels. The structure is shown in Fig. 17, which is denoted as ”ResnetBlock (Attention)” (RBA) in the figure.
This attention mechanism incorporates a learnable 2D Rotary Position Embedding (RoPE), which has demonstrated a remarkable ability to handle multi-scale inputs and extrapolate across resolutions [heo2024RoPE2d]. By applying rotational transformations to the query and key vectors based on their spatial coordinates, 2D RoPE effectively encodes relative positional information directly into the attention mechanism. This allows the model to capture complex, long-range dependencies between spin sites across the entire lattice, a crucial capability for modeling systems at or near a phase transition where correlation lengths diverge. The frequencies of the rotary embeddings are treated as learnable parameters, enabling the model to adapt the positional encoding to the specific physics of the system.
D.2 Optimization of Neural Network Architecture
The entire model is optimized using ”torch.compile” to enhance computational efficiency [pytorch]. For certain blocks (”ResnetBlock” and ”ResnetBlock (Attention)” in Fig. 17) in model is optimized using ”epilogue fusion” and ”max autotune” supported by PyTorch [pytorch] to perform operator fusion automatically. Please refer to the code in the following link for the detailed structure of the neural network. (https://github.com/ToelUl/Flow-U-Net)
Figure 17: The core Building Blocks in U-Net. The left-hand side is ”ResnetBlock (Attention)” (RBA), which is used in the network’s deeper, lower resolution levels. ”ResnetBlock” (RB) and ”ResnetBlock (Attention)” (RBA) share the same structure except for the attention block. Please refer to the code in the following link for the detailed structure of the neural network. (https://github.com/ToelUl/Flow-U-Net)
Appendix E Comparison between U-Net and DiT Architectures
To show that the inductive biases embedded in CNN architectures of U-Net are the key point for training on data of a fixed system size and extrapolating to other sizes, we compare the results using another robust neural network architecture: the Diffusion Transformer [peebles2023DiT].
Vision Transformers [vit] and, by extension, the Diffusion Transformer (DiT) [peebles2023DiT] variants recently adopted for generative modeling, rely on minimal inductive biases. Transformer’s core is the self-attention mechanism [vaswani2023attention_need]. This mechanism empowers them with global receptive fields and permutation equivariance, enabling each image patch (or token) to interact with all others regardless of spatial proximity. The only explicit inductive bias is introduced through the positional encoding scheme, which provides spatial context but does not enforce strict locality. As a consequence, Transformers excel when long-range correlations or nonlocal dependencies dominate the target distribution, conditions frequently encountered in natural language, graphs, and physical systems near criticality where correlation lengths diverge. But in our case, the underlying data-generating process is governed by a physical Hamiltonian with strictly local interactions.
We conducted a comparative experiment under identical training conditions to validate this hypothesis. Both the U-Net and DiT models were trained on the same dataset from MCMC simulations of a $32\times 32$ system. Notably, to specifically accommodate the potentially higher data requirements of the larger DiT model, we increased the number of training samples to $2500$ configurations per temperature point (up from 1500 used in the system size extrapolation experiment, see Sec. V). Both models were trained for the same number of epochs with an identical optimizer and learning rate schedule, ensuring a controlled comparison. Both models ensure that the training has converged and the training loss curve is shown in the Fig. 18.
Figure 18: Training loss curves of U-Net and DiT, where each point represents the average loss per epoch. The two models exhibit similar convergence behavior, but U-Net performs better. Inference sampling results for the two models are shown in Figs. 19 and  20.
A potential concern in this comparison is the significant disparity in model size (Table 1), where the DiT model has nearly six times more parameters than the U-Net. One might argue that the DiT’s poorer performance could stem from a larger, more complex model that is inherently more difficult to train, rather than from a fundamental misalignment of its inductive biases.
Table 1: Comparison of U-Net and DiT model architectures. Due to its structural limitations, the DiT model is significantly larger regarding parameters and computational cost (floating point operations, FLOPs). All training and inference conditions were held constant between the two models.
Metric U-Net DiT
Parameters 214.65 K 1262.98 K
FLOPs 0.9776 G 6.4847 G
Memory (params) 0.82 MB 4.82 MB
However, we argue that this disparity in parameters is not a confounding factor but a direct consequence of the architectural differences and is, in fact, central to our conclusion. The DiT architecture, lacking a built-in locality bias, requires a much larger parameter space to learn local physical relations from data—a task for which the U-Net is purpose-built via its convolutional structure. The DiT’s substantial parameter count is a necessity from its lack of correct inductive bias. Therefore, its struggle to train effectively under these conditions is not an artifact of an unfair comparison but rather a direct demonstration of our central thesis: for physical systems governed by local interactions, having the correct architectural inductive bias is far more critical than raw model capacity. The training challenge itself highlights the inherent efficiency and superiority of the U-Net’s architecture for this class of problems.
The DiT architecture, lacking strong locality bias, must deduce the principle of locality entirely from the training data. Its self-attention mechanism initially treats all pairs of sites as potentially relevant, placing a significant learning burden on the model to discover that the system’s dynamics are overwhelmingly dominated by nearest-neighbor interactions. This leads to a critical challenge in generalization, as observed in our Flow Matching experiments (see Figs. 19 and 20). The DiT model tends to learn correlations and patterns that are specific to the fixed system size seen during training, rather than the universal, size-invariant rule of local physical law. When extrapolating to larger, unseen system sizes ($64\times 64$), the global attention patterns learned in the smaller system are no longer directly applicable, leading to a significant performance degradation.
Conversely, the powerful inductive biases of a CNN-based architecture in the U-Net align perfectly with the fundamental structure of the local Hamiltonian. The convolutional kernel intrinsically learns a local rule that is, by its very nature, independent of the total system size. This inherent understanding of locality and translation equivariance means the learned physical principles are more robust and readily generalizable. The network does not need to waste capacity learning the concept of locality; it can instead focus on modeling the specific nature of these local interactions. This alignment explains why the U-Net architecture demonstrates superior extrapolation capabilities, successfully generalizing to different system sizes after being trained on only one.
This observation is consistent with a large body of literature in computer vision comparing Transformers and CNNs. The seminal work on Vision Transformers by Dosovitskiy et al. [vit] explicitly states that Transformers lack the inductive biases of locality and translation equivariance, making them heavily reliant on massive-scale pre-training to generalize effectively. Subsequent analyses, such as the work by Raghu et al. [vit_vs_cnn], have provided further evidence, showing that ViTs must learn to represent local information from scratch, a task for which CNNs are architecturally predisposed. Our comparative experiment with the DiT model, therefore, serves as a direct demonstration of this principle within a physical simulation context: the model with the correct inductive bias (U-Net) not only performs better but does so with a fraction of the parameters, because its architecture is already aligned with the fundamental, local nature of the physical system.
Figure 19: This figure shows the performance of the U-Net and DiT trained and sampled at the same size ($32\times 32$). The MCMC (MC) data used for training and comparison have a temperature interval 0.05. The sampling temperature interval of U-Net and DiT after training with Flow Matching is 0.01. The figure clearly shows that even when sampling at the training size, DiT, which lacks the correct inductive bias, cannot learn the correct physical behavior with limited training data like the U-Net.
Figure 20: This figure shows the performance of U-Net and DiT trained at size $32\times 32$ and sampled at a larger size $64\times 64$. The MCMC data used for training at size $32\times 32$ had a temperature interval of 0.05. U-Net and DiT, after training with Flow Matching, had a sample temperature interval of 0.01. The MCMC (MC) data used for comparison at the larger size $64\times 64$ in the figure had a temperature interval of 0.01. The figure shows that DiT, which lacks the correct inductive bias, has worse results under size extrapolation (compared to Figs. 19) and does not exhibit correct finite-size scaling behavior.
Appendix F GPU-accelerated Monte Carlo Method
We employ a highly optimized MCMC simulation framework implemented in PyTorch [pytorch]. While the Wolff [PhysRevLett_Wolff] or Swendsen–Wang [PhysRevLett.58.86] cluster algorithm is known to be more efficient in reducing autocorrelation steps for the 2D XY model, its sequential cluster-building process is not amenable to efficient GPU parallelization. Our choice of the Metropolis-Hastings algorithm with a checkerboard update scheme is motivated by its exceptional performance on massively parallel architectures, making it a powerful and relevant baseline for modern hardware. The full implementation of our algorithm is available on GitHub: (https://github.com/ToelUl/Thermal-Phase-Transition-MC-GPU-Simulation). The algorithm is designed to leverage the massively parallel architecture of modern Graphics Processing Units (GPUs), enabling large-scale simulations with high statistical precision. The core design incorporates three key parallelization strategies: a checkerboard update scheme for intra-lattice parallelism, the simulation of multiple independent Markov chains, and the replica-exchange method (also known as Parallel Tempering [Hukushima_1996]) across a range of temperatures.
F.1 Parallel Architecture and Data Layout
The fundamental data structure for the spin configurations is a four-dimensional tensor of shape $B\times C\times L\times L$, where:
  • $B$ is the batch size, representing the number of different temperatures simulated simultaneously.
  • $C$ is the number of independent Monte Carlo chains for each temperature.
  • $L$ is the linear size of the square lattice.
This layout allows for all temperatures and chains to be processed in parallel. Operations are vectorized across the first two dimensions ($B$ and $C$), which maps naturally to the single instruction, multiple data (SIMD) paradigm of GPU computing. The use of $C$ independent chains per temperature not only accelerates the accumulation of statistics but also enhances the statistical independence of the collected samples, providing a more robust ensemble average.
F.2 Checkerboard Metropolis-Hastings Update
A standard single-spin flip Metropolis-Hastings algorithm introduces data dependencies, as the update of a spin requires knowledge of its neighbors, precluding simultaneous updates of adjacent spins. To overcome this limitation and maximize parallel efficiency, we implement a checkerboard update scheme [Weigel_2011]. The lattice is decomposed into two independent sub-lattices, ’A’ and ’B’ (analogous to the black and white squares of a checkerboard).
A single Monte Carlo sweep consists of two sequential steps:
  1. 1.
    All spins on sub-lattice A are updated simultaneously. Since no spin on sub-lattice A is a nearest neighbor to any other spin on A, these updates are fully independent and can be performed in a single, vectorized operation on the GPU.
  2. 2.
    All spins on sub-lattice B are then updated simultaneously in the same manner.
For the XY model, where the Hamiltonian is given by:
$H=-J\sum_{\langle i,j\rangle}\cos(\theta_{i}-\theta_{j})$ (34)
the local update involves proposing a new angle $\theta^{\prime}_{i}=\theta_{i}+\delta\theta$ for a spin $i$, where $\delta\theta$ is a random perturbation. The change in energy, $\Delta E$, is computed based only on the local spin and its neighbors. The proposed move is then accepted with the probability:
$P_{\mathrm{acc}}(\theta_{i}\to\theta^{\prime}_{i})=\min\left(1,e^{-\Delta E/T}\right)$ (35)
where TT is the temperature of the specific replica. This entire process, from proposing moves to accepting them, is executed in parallel for all spins on the active sub-lattice across all chains and temperatures.
F.3 Parallel Tempering (Replica-Exchange)
To mitigate critical slowing down near phase transitions and to ensure efficient exploration of the configuration space, our framework implements the Parallel Tempering (or replica-exchange) algorithm [Hukushima_1996]. This method involves running simulations at multiple temperatures $T_{1},T_{2},\dots,T_{B}$ in parallel. Periodically, an exchange of the entire spin configurations between two adjacent replicas, say at temperatures $T_{i}$ and $T_{i+1}$ with energies $E_{i}$ and $E_{i+1}$, is proposed.
The exchange is accepted with a probability that satisfies the detailed balance condition:
$P(\mathrm{swap})=\min\left(1,\exp\left[\left(\frac{1}{T_{i}}-\frac{1}{T_{i+1}}\right)(E_{i}-E_{i+1})\right]\right)$ (36)
This allows configurations from high-temperature replicas, which explore the energy landscape more broadly, to diffuse to lower temperatures, preventing them from getting trapped in local energy minima. To ensure detailed balance over the entire system, exchange attempts are alternated between pairs $(T_{i},T_{i+1})$ with even $i$ and those with odd $i$. This technique significantly improves the convergence and ergodicity of the simulation, especially for complex systems.
By combining these three levels of parallelism, the algorithm achieves a substantial reduction in simulation time while generating statistically robust and independent samples across a wide range of temperatures.
Appendix G Information about Computing Resources.
All numerical experiments, including model training, inference, and Monte Carlo simulations, were conducted on a workstation with the following specifications.
  • CPU: AMD Ryzen 9 9950X 16-Core Processor
  • GPU: 1x NVIDIA GeForce RTX 4080 (16 GB VRAM)
  • Operating System: Windows Subsystem for Linux 2 (WSL2)
  • Core Libraries [pytorch, cudnn, lipman2024flow_matching_guide_code]:
    • PyTorch: 2.7.1 (+cu128)
    • CUDA Toolkit (used by PyTorch): 12.8
    • cuDNN: 9.7.1
    • flow_matching: 1.0.10